In Huffman's coding, if the symbol A has probability 0.3, B has probability 0.15, C has probability 0.1, D has probability 0.25 and E has probability 0.2,…
In Huffman's coding, if the symbol A has probability 0.3, B has probability 0.15, C has probability 0.1, D has probability 0.25 and E has probability 0.2, then the minimum number of bits required to represent all the symbols together is
Show answerHide answer
Correct answer: (C) 12
Explanation
Huffman merges: 0.1 + 0.15 = 0.25, then 0.2 + 0.25 = 0.45, then 0.25 + 0.3 = 0.55, then 1. The code lengths are A 2, D 2, E 2, B 3, C 3, totalling 2 + 2 + 2 + 3 + 3 = 12 bits.